Math 210A Homework Assignment (11/04 – 11/18)

Assignment will be collected on 11/28 the last Monday in November during the class.

Key notions introduced this long two weeks period: (1) Isolated singularities, removable singularity, pole, essential sigularity;

(2) Laurent series;

(3) Casorati-Weierstrass Theorem;

(4) Residues;

(5) Argument principle, Rouche’s Theorem.

Homework problems:

§5.1 1, 4. §5.2. 1(a), (b), (d). §5.3. 1, 2.